OpenStudy (eunice_o):

Divide. 16x^6 -18x^3/2x^3

OpenStudy (eunice_o):

@RhondaSommer

RhondaSommer (rhondasommer):

okay so \[16x^6 -18x^3/2x^3\]do 16/2=? and -18/2=?

OpenStudy (eunice_o):

8 and -9

RhondaSommer (rhondasommer):

okay so now\[\frac{x^6}{x^3}\]is basically \[x^{6-3}\]so what is 6-3?

OpenStudy (eunice_o):

3

RhondaSommer (rhondasommer):

yup! okay so now we have\[8x^3-9\]now\[\frac{x^3}{x^3}\]is basically\[x^{3-3}\]so what is \[3-3\]

OpenStudy (eunice_o):

0

RhondaSommer (rhondasommer):

yup. so we dont put that one in :D so your answer is...

OpenStudy (eunice_o):

8x^3-9

RhondaSommer (rhondasommer):

yup :D

OpenStudy (eunice_o):

so basically we used the denominator to get the answers from the numerator ?

RhondaSommer (rhondasommer):

yes, you subtracted the numerator from the denominator (only the exponents)

RhondaSommer (rhondasommer):

had the number been a negative you would have place everything but that exponent on top and make it a positive (say it had been \[4x^{3-6}y\] then it would be \[4\frac{x^3}{y}\]

OpenStudy (mathstudent55):

\(\Large \dfrac{16x^6 -18x^3}{2x^3} =\) \(\Large = \dfrac{16x^6}{2x^3} - \dfrac{18x^3}{2x^3} \) \(\Large = \dfrac{16}{2}x^{6 - 3} - \dfrac{18}{2} x^{3 - 3} \) \(\Large = 8x^3 - 9x^0\) \(\Large =8x^3 - 9\)

OpenStudy (mathstudent55):

\(\Large 4x^{3 - 6}y = 4x^{-3}y = \dfrac{4y}{x^3}\)

RhondaSommer (rhondasommer):

we did that